Determine the Keq for the following equilibrium:

2CH4(g) ⇌ C2H2(g) + 3H2(g)

(CH4 = 2.5 x 10-3 M ; C2H2 = 0.01375 ; H2 = 0.04125)


1
Expert's answer
2022-05-23T04:16:03-0400

"K_{eq}=\frac{[H_2]^3[C_2H_2]}{[CH_4]^2}=\frac{0.04125^3(0.01375)}{(2.5x10^{-3})^2}=0.154"